# Quantum Signatures of Strange Attractors

### v2

@@ I Introduction
In this work, we extend the CK approach to driven, dissipative mechanical systems and present, for the first time, phase‐space portraits of its strange attractor in terms of the Husimi distributions[8]. We demonstrate our approach on the Duffing equation[9]. By tracing their evolution under nonlinear driving and damping, we can visualize how dissipation sculpts the Husimi function into stretched, filamentary structures that mirror classical chaotic saddles. The Husimi smoothing[8] is then applied to reveal the gradual loss of coherence along unstable manifolds. In doing so, we provide a complementary, geometric viewpoint to more familiar Lindblad- and Langevin-based analyses[10, 11, 12], highlighting the distinct signatures of pure dissipation in the quantum-chaotic phase space.
@@ III Semiclassification
$\frac{d}{dt}\langle\psi|{\bf x}|\psi\rangle$ $=$ $\frac{1}{m}\langle\psi|{\bf P}|\psi\rangle,$
$\frac{d}{dt}\langle\psi|{\bf P}|\psi\rangle$ $=$ $-\delta(t)\langle\psi|{\bf P}|\psi\rangle-\langle\psi|\nabla V({\bf x},t)|\psi\rangle,$ (3)
@@ VI Chaos and Strange Attractors in the Duffing Equation
One remarkable feature of the Duffing oscillator is the period-doubling cascade to chaos. As the forcing amplitude $\gamma$ or frequency $\omega$ is varied, the system undergoes a sequence of period-doubling bifurcations, wherein stable periodic orbits repeatedly bifurcate, doubling their period at each bifurcation. This cascade continues until the system dynamics become chaotic. The presence of chaotic behavior can be quantified through Lyapunov exponents, with positive exponents indicating sensitive dependence on initial conditions, a hallmark of chaos [17].
@@ VII Numerical solution
$\left(e^{-i\hat{T}_{mid}\delta t}\psi\right)(x)=\mathcal{F}^{-1}\left\{\exp\left(-i\frac{a(t_{mid})\delta t}{2}k^{2}\right)\mathcal{F}\{\psi(k)\}\right\}$

### v3

@@ I Introduction
In this work, we extend the CK approach to driven, dissipative mechanical systems and present, for the first time, phase‐space portraits of its strange attractor in terms of the Husimi distributions[8]. We demonstrate our approach on the Duffing equation[9, 10]. By tracing their evolution under nonlinear driving and damping, we can visualize how dissipation sculpts the Husimi function into stretched, filamentary structures that mirror classical chaotic saddles. The Husimi smoothing[8] is then applied to reveal the gradual loss of coherence along unstable manifolds. In doing so, we provide a complementary, geometric viewpoint to more familiar Lindblad- and Langevin-based analyses[11, 12, 13, 14], and the non-Hermitian approach of Iomin and Zaslavsky[15], highlighting the distinct signatures of pure dissipation in the quantum-chaotic phase space.
The central aim of this study is to characterize the phase-space morphology of dissipative quantum chaos, specifically addressing how the fractal geometry of classical strange attractors survives the smoothing effects of the uncertainty principle. By utilizing the Caldirola-Kanai Hamiltonian, we circumvent the phenomenological decoherence inherent in Lindblad formalisms to isolate the deterministic role of friction in quantum state evolution. The significance of this investigation for nonlinear dynamics lies in establishing that dissipation acts as a requisite localizing force, collapsing the wavefunction onto the classical invariant set. Our analysis of the out-of-time-ordered correlator (OTOC) provides a quantitative criterion for the quantum-classical transition, demonstrating that the recovery of the classical Lyapunov exponent is strictly dependent on the dissipation rate driving the system toward the semiclassical limit.
@@ III Semiclassification
$\frac{d}{dt}\langle\psi|{\bf x}|\psi\rangle$ $=$ $\frac{1}{m}\langle\psi|{\bf P}|\psi\rangle,$ (3)
$\frac{d}{dt}\langle\psi|{\bf P}|\psi\rangle$ $=$ $-\delta(t)\langle\psi|{\bf P}|\psi\rangle-\langle\psi|\nabla V({\bf x},t)|\psi\rangle,$
@@ VI Chaos and Strange Attractors in the Duffing Equation
One remarkable feature of the Duffing oscillator is the period-doubling cascade to chaos[20]. As the forcing amplitude $\gamma$ or frequency $\omega$ is varied, the system undergoes a sequence of period-doubling bifurcations, wherein stable periodic orbits repeatedly bifurcate, doubling their period at each bifurcation. This cascade continues until the system dynamics become chaotic. The presence of chaotic behavior can be quantified through Lyapunov exponents, with positive exponents indicating sensitive dependence on initial conditions, a hallmark of chaos [21].
@@ VII Numerical solution
$\left(e^{-i\hat{T}_{mid}\Delta t}\psi\right)(x)=\mathcal{F}^{-1}\left\{\exp\left(-i\frac{a(t_{mid})\Delta t}{2}k^{2}\right)\mathcal{F}\{\psi(x)\}\right\}$